OpenStudy (anonymous):

Using the change of base formula, evaluate log312 (round to the nearest hundredth). 2.26 0.48 4 1.08

OpenStudy (campbell_st):

well change of base is simply changing to a base 10 or base e log \[\log_{a}(b) = \frac{\ln(b)}{\ln(a)}\] then evaluating in your question a = 3 and b = 12 just substitute and evaluate...

OpenStudy (anonymous):

Thank you.